# main_arXiv.bib

### v1

@@ 2.3 Value Function and Q-function
We follow the approach of [asienkiewicz2017note] although as they prove that among all history dependent policies there exists an optimal stationary policy we will reduce the notation by considering only stationary policies from the outset.
@@
Let $v\in\mathbb{R}^{S}$ and $\pi:\mathcal{S}\rightarrow\mathcal{A}$ be a stationary deterministic policy. We then define $\rho_{s}^{a}:\mathbb{R}^{S}\rightarrow\mathbb{R}$ by
$\rho_{s}^{a}(v)=-\frac{1}{\beta}\log\bigg{(}\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta v(s^{\prime})}]\bigg{)}$ (5)
[…]
@@
The problem of the agent is then for all initial states $s\in\mathcal{S}$ to find a policy $\pi^{*}$ that solves $J(s,\pi^{*})=\sup_{\pi}J(s,\pi)$. In [asienkiewicz2017note] the authors consider a more general framework than is not restricted to finite MDPs or stationary policies but they prove that under some conditions (that are trivially fulfilled in the case of finite MDPs) there exists a stationary policy $\pi^{*}$ that maximizes the state-value function for all states $s\in\mathcal{S}$ simultaneously. The authors only prove this in the risk-averse case $\beta>0$ but the proof carries over verbatim in the case of $\beta<0$. They also show that any optimal policy $\pi^{*}$ that solves the agent satisfies the optimality equation
@@
$V^{*}(s)=\max_{a\in\mathcal{A}}\bigg{(}R(s,a)-\frac{\gamma}{\beta}\log\bigg{(}\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta V^{*}(s^{\prime})}]\bigg{)}\bigg{)}$ (6)
where $V^{*}$ is the optimal $V$-function.
@@
Also for any stationary deterministic policy $\pi$ we have the Bellman recursion:
$V^{\pi}(s)=R(s,\pi(s))-\frac{\gamma}{\beta}\log\bigg{(}\mathbb{E}_{s^{\prime}\sim P(\cdot|s,\pi(s))}[e^{-\beta V^{\pi}(s^{\prime})}]\bigg{)}$ (7)
Since we are not interested only in the planning problem but that of learning, we also introduce the state-action value function $Q$. The approach is very similar to that of the value-function. Given $\pi$, we define the operator $L_{\pi}:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S\times A}$ as follows: for all $v:\mathcal{S}\to\mathbb{R}$,
$L_{\pi}(v)(s,a)=R(s,\pi(s))+\gamma\rho_{s}^{\pi}(v)$ (8)
We define the operator $L:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S\times A}$ for all $v:\mathcal{S}\rightarrow\mathbb{R}$ as:
$L(v)(s,a)=R(s,a)+\gamma\rho_{s}^{a}(v)$
@@
We define the $N$-step total discounted utility of the state-action pair $(s,a)$ under $\pi$ as $L_{N}(s,a,\pi):=(L\circ J_{\pi}^{N-1}(\boldsymbol{0}))(s,a)$ and the limit is denoted $Q^{\pi}(s,a)$: $Q^{\pi}(s,a)=\lim_{N\rightarrow\infty}L_{N}(s,a,\pi)$. Although the authors do not consider state-action value functions in their paper, repeating the arguments of [asienkiewicz2017note] it suffices to consider stationary policies when wanting to solve $\max_{\pi}Q^{\pi}(s,a)$ for all $(s,a)$ and that the solution $Q^{*}$ solves the optimality equation
$Q^{*}(s,a)=R(s,a)-\frac{\gamma}{\beta}\log\bigg{(}\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta\max_{a}Q^{*}(s,a)}]\bigg{)}$ (9)
@@
Similarly it is clear that the $Q$-functions satisfy the Bellman recursion relations
$Q^{\pi}(s,a)=R(s,a)-\frac{\gamma}{\beta}\log\big{(}\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta V^{\pi}(s^{\prime})}]\big{)}$ (10)
@@ 2.4 Learning Performance
We consider two types of RL algorithms $\mathcal{U}$, namely those that output a $Q$-function $Q_{T}^{\mathcal{U}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ and those that output a policy $\pi_{T}^{\mathcal{U}}:\mathcal{S}\rightarrow\mathcal{A}$ using $T$ transition samples. Note that any algorithm that outputs a $Q$-function also outputs a policy, namely the one obtained by acting greedy with respect to the $Q$-function. There are also ways to obtain a $Q$-function from a policy. There is however no canonical way to do this as the algorithm cannot simply output $Q^{\pi^{\mathcal{U}}}$ since the algorithm does not have access to the true MDP. The way we evaluate the quality of an algorithm that outputs a $Q$-function is by $\|Q^{*}-Q_{T}^{\mathcal{U}}\|$. For algorithms that instead outputs a policy we evaluate the policy in terms of $\|V^{*}-V^{\pi_{T}^{\mathcal{U}}}\|$. Often we will suppress $T$ from the notation.
@@
We say that a algorithm $\mathcal{U}$ that outputs a $Q$-function $Q^{\mathcal{U}}$ is $(\varepsilon,\delta)$-correct on a set of MDPs $\mathbb{M}$ if $\mathbb{P}(\|Q^{*}-Q^{\mathcal{U}}\|\leq\epsilon)\geq 1-\delta$ for all $M\in\mathbb{M}$. Similarly we say that an algorithm $\mathcal{U}$ that outputs a policy $\pi^{\mathcal{U}}$ is is $(\varepsilon,\delta)$-correct on a set of MDPs $\mathbb{M}$ if $\mathbb{P}(\|V^{*}-V^{\mathcal{\pi^{\mathcal{U}}}}\|\leq\epsilon)\geq 1-\delta$ for all $M\in\mathbb{M}$.
@@ 3 Model-Based Risk-Sensitive Q-Value Iteration
In this section we describe the model-based value iteration algorithm which aims at finding the optimal $Q$-function $Q^{*}$. We then give an upper bound on the total number of calls to the generative model needed in order for this algorithm to be $(\varepsilon,\delta)-$correct. The model based approach is based on working on a model MDP which can disagree with the true MDP because it does not use the true transition probabilities but an estimate of the transition functions obtained from $n$ calls to each of the state-action pairs in $\mathcal{S}\times\mathcal{A}$ as described in the algorithm below.
@@ 12            $\forall s\in\mathcal{S}:\widehat{P}(s,z)=\frac{m(s,z)}{n}$
The model-based approach we describe is general in the sense that any oracle that for any $\varepsilon>0$ can find an $\varepsilon$-optimal policy can be used. We prove the existence of one such oracle in the form of a $Q$-value iteration very like the one from the classical risk-neutral setting. The proof is very similar to Part (a) in Theorem 3.1 in [asienkiewicz2017note] but is nevertheless provided in Appendix B.
@@
$(\mathcal{T}f)(s,a)$ $=R(s,a)+\frac{-\gamma}{\beta}\log\bigg{(}\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta\max_{a}f(s^{\prime},a)}\bigg{)}$
$(\mathcal{T}^{\pi}f)(s,a)$ $=R(s,a)+\frac{-\gamma}{\beta}\log\bigg{(}\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta f(s^{\prime},\pi(s^{\prime}))}\bigg{)}$
The operators $\mathcal{T},\mathcal{T}^{\pi}$ are $\gamma$-contractions with respect to the max-norm, i.e. for value-functions $f_{1}$ and $f_{2}$ it holds that $\|\mathcal{T}f_{1}-\mathcal{T}f_{2}\|\leq\gamma\|f_{1}-f_{2}\|$ and $\|\mathcal{T}^{\pi}f_{1}-\mathcal{T}^{\pi}f_{2}\|\leq\gamma\|f_{1}-f_{2}\|$.
@@ 8                   $Q_{j+1}(s,a)=\mathcal{T}Q_{j}(s,a)$
Algorithm 2 RS-QVI$(M,k)$
The next lemma shows that if we choose $k$ large enough in the RS-QVI algorithm, we can obtain $Q_{k}$ and $V^{\pi_{k}}$ that are as close to $Q^{*}$ and $V^{*}$ as we desire. The proof is postponed to Appendix B.
@@
Fix $\varepsilon>0$. Then there exists some $k(\varepsilon)$ such that if the number of iterations in RS-QVI uses more than $k(\varepsilon)$ iterations then, the output of the Algorithm 2 (RS-QVI) satisfies $\|Q_{k}-Q^{*}\|<\varepsilon$ and $\|V^{\pi_{k}}-V^{*}\|<\varepsilon$
@@ 3.1 Analysis of MB-RS-QVI
With RS-QVI in place we need a set of lemmas for the analysis of the sample-complexity of the model based RS-QVI algorithm.
An important result for the analysis is a risk-sensitive version of the simulation lemma [kearns2002near, strehl_analysis_2008], which describes how different two $Q$-functions for the same policy are in two different MDPs that differ only slightly in their rewards and transition functions. The proof is postponed to Appendix B.
@@
Consider two MDPs $M_{1}=(S,A,P_{1},R,\gamma,\beta)$ and $M_{2}=(S,A,P_{2},R,\gamma,\beta)$ differing only in their transition functions. Fix a stationary policy $\pi$, and let $Q^{\pi}_{1}$ and $Q^{\pi}_{2}$ be respective the corresponding $Q$-functions of $\pi$ in $M_{1}$ and $M_{2}$. If $\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\|P_{1}(\cdot|s,a)-P_{2}(\cdot|s,a)\|_{1}\leq\tau$. Then, it holds that
$\|V_{1}^{\pi}-V_{2}^{\pi}\|\leq\|Q^{\pi}_{1}-Q^{\pi}_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\tau$
It then follows that if for some $\varepsilon>0$ it holds that $\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\|P_{1}(\cdot|s,a)-P_{2}(\cdot|s,a)\|_{1}\leq\varepsilon\frac{1-\gamma}{\gamma}|\beta|e^{-|\beta|\frac{1}{1-\gamma}}$, then $\|Q_{1}^{\pi}-Q_{2}^{\pi}\|\leq\varepsilon$. Using the decompositions from Lemma 5 we get for any $s\in\mathcal{S}$ that
$V^{\pi_{k}}(s)$ $\geq V^{*}(s)-\|V^{\pi_{k}}-\widehat{V}^{\pi_{k}}\|-\|\widehat{V}^{\pi^{*}}-V^{*}\|-\|\widehat{V}^{\pi_{k}}-\widehat{V}^{*}\|$ (11)
@@
Lemma 4.
Fix $\tau>0$. If every state-action pair has been tried at least $m=8\frac{S+\log(\frac{SA}{\delta})}{\tau^{2}}$ times, then it holds that $\max_{(s,a)}\|P(\cdot|s,a)-\hat{P}(\cdot|s,a)\|_{1}\leq\tau$ with probability at least $1-\delta$.
@@
Using the Weissman inequality [weissman2003inequalities], any confidence interval for a state-action pair that have been tried $m$ times have size $2\sqrt{2[\log(2^{S}-2)-\log(\delta_{P})]/m}$. Setting this size to be smaller than $\tau$ and solving for $m$, we find that
$m\geq\frac{8}{\tau^{2}}\Big{(}\log(2^{S}-2)-\log(\delta_{P})\Big{)}\,.$ (13)
Noting that $8\big{(}S+\log(1/\delta_{P})\big{)}/\tau^{2}\geq 8[\log(2^{S}-2)-\log(\delta_{P})]/\tau^{2}$ and substituting $\delta_{P}=\frac{\delta}{SA}$, the result follows by the union bound since each of the $SA$ confidence balls contain $P(\cdot|s,a)$ with probability at least $1-\delta$. ∎
@@
There exist a universal constant $c$ such that for any $\varepsilon>0,\delta\in(0,1)$ and any MDP $M$ with $S$ states and $A$ actions if the learner is allowed to make
@@
calls to the generative model then $\mathbb{P}(\|Q^{*}-Q_{k}\|\leq\varepsilon)\geq 1-\delta$
@@
For any $\varepsilon>0$, we can get $\|Q_{k}-\widehat{Q}^{*}\|<\varepsilon/2$ using enough iterations of the optimization oracle by Lemma 2. The term $\|\widehat{Q}^{\pi^{*}}-Q^{*}\|$ can also be made smaller than $\varepsilon/2$ by the simulation lemma if $\max_{(s,a)}\|P(\cdot|s,a)-\widehat{P}(\cdot|s,a)\|_{1}<\tau$ where $\tau=\frac{\varepsilon}{2}\bigg{[}\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\bigg{]}^{-1}$ which can be ensured with probability larger than $1-\delta$ by picking $m=8\frac{S+\log(\frac{SA}{\delta})}{\tau^{2}}$. Using that the total calls to the generative model is $T=SAm$ and substituting in the value for $\tau$ we can ensure for all $(s,a)$ that $Q^{\pi}(s,a)>Q^{*}(s,a)-\varepsilon$ with probability larger than $1-\delta$ by using a total number of samples
$T=32\bigg{[}S+\log\bigg{(}\frac{SA}{\delta}\bigg{)}\bigg{]}\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{2|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}$ (15)
Theorem 2.
There exist a universal constant $c$ such that for any $\varepsilon>0,\delta\in(0,1)$ and any MDP $M$ with $S$ states and $A$ actions if the learner is allowed to make
calls to the generative model then $\mathbb{P}(\|V^{*}-V^{\pi_{k}}\|\leq\varepsilon)\geq 1-\delta$.
Proof.
By Lemma 2 we can pick $k$ large enough so that $\|\hat{V}^{\pi_{k}}-\hat{V}^{*}\|\leq\varepsilon/3$. The two terms $\|V^{\pi_{k}}-\widehat{V}^{\pi_{k}}\|$ and $\|\hat{V}^{\pi^{*}}-V^{*}\|$ can simultaneously be made smaller than $\varepsilon/3$ by sampling each state-action pair $m=8\frac{S+\log(\frac{SA}{\delta})}{\tau^{2}}$ times where $\tau=\frac{\varepsilon}{3}\bigg{[}\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\bigg{]}^{-1}$ such that if we sample
$T=72\bigg{[}S+\log\bigg{(}\frac{SA}{\delta}\bigg{)}\bigg{]}\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{2|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}$ (17)

### v2

@@ 2.3 Value Function and Q-function
We follow the approach of asienkiewicz2017note and bauerle_markov_2024 but since none of their cases include our $\beta>0$ case and also only cover $V$-functions, we give in Section C the full setup with history-dependent policies as well as a full definition of the $V$ and $Q$-functions, and prove existence of a stationary optimal policy and show that the value functions of this policy satisfy a Bellman optimality equation and similarly that value functions of any policy satisfy a Bellman recursion relation. We give an outline here that only deals with stationary policies, which is justified by the results of Section C.
@@
Let $v\in\mathbb{R}^{S}$ and $\pi:\mathcal{S}\rightarrow\mathcal{A}$ be a stationary deterministic policy. We then define $\rho_{s,a}:\mathbb{R}^{S}\rightarrow\mathbb{R}$ by
$\rho_{s,a}(v)=-\frac{1}{\beta}\log\bigg(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta v(s^{\prime})}]\bigg)$ (5)
and slighltly abusing the notation, we write $\rho_{s,\pi}$ when $a=\pi(s)$, that is $\rho_{s,\pi}:=\rho_{s,\pi(s)}$. We then define the operator $J_{\pi}:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S}$ given by $J_{\pi}(v)(s)=r(s,\pi(s))+\gamma\rho_{s,\pi}(v)$. The $N$-step total discounted utility $J_{N}(s,\pi)$ is defined as applying $J_{\pi}$ recursively $N$-times to the 0-map, that is $J_{N}(s,\pi):=J_{\pi}^{N}(\boldsymbol{0})(s)$. Note that the outer-most iteration corresponds to the immediate time-step. By properties (1)-(2) of $\rho$, it follows that $J_{N}(s,\pi)$ is increasing in $N$ and that $J_{N}(s,\pi)\leq\frac{1}{1-\gamma}$, so that the limit $N\rightarrow\infty$ exists. This limit is considered the value of state $s$ under the policy $\pi$ of the agent: $V^{\pi}(s)=\lim_{N\rightarrow\infty}J_{N}(s,\pi)$.
@@
The problem of the agent is then for all initial states $s\in\mathcal{S}$ to find a policy $\pi^{*}$ that solves $J(s,\pi^{*})=\sup_{\pi}J(s,\pi)$. In asienkiewicz2017note , the authors consider a more general framework that is not restricted to finite MDPs or stationary policies; they prove that under some conditions —that are trivially fulfilled in the case of finite MDPs— there exists a stationary policy $\pi^{*}$ that maximizes the state-value function for all states $s\in\mathcal{S}$ simultaneously. We show that any optimal policy $\pi^{*}$ that solves the agent satisfies the optimality equation:
@@
$V^{*}(s)=\max_{a\in\mathcal{A}}\bigg(R(s,a)-\frac{\gamma}{\beta}\log\bigg(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta V^{*}(s^{\prime})}]\bigg)\bigg)\,$, (6)
@@
where $V^{*}$ is the optimal $V$-function. Also for any stationary deterministic policy $\pi$, we have the Bellman recursion:
$V^{\pi}(s)=R(s,\pi(s))-\frac{\gamma}{\beta}\log\bigg(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,\pi(s))}[e^{-\beta V^{\pi}(s^{\prime})}]\bigg)\,.$ (7)
@@
Since we are not interested only in the planning problem but also in learning, we also introduce the state-action value function $Q$. The approach is very similar to that of the value-function. Given $\pi$, we define the operator $L_{\pi}:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S\times A}$ as follows: for all $v:\mathcal{S}\to\mathbb{R}$, $L_{\pi}(v)(s,a)=R(s,\pi(s))+\gamma\rho_{s,\pi}(v)\,.$ We define the operator $L:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S\times A}$ for all $v:\mathcal{S}\rightarrow\mathbb{R}$ as: $L(v)(s,a)=R(s,a)+\gamma\rho_{s,a}(v)\,.$ We define the $N$-step total discounted utility of the state-action pair $(s,a)$ under $\pi$ as $L_{N}(s,a,\pi):=(L\circ J_{\pi}^{N-1}(\boldsymbol{0}))(s,a)$ and the limit is denoted $Q^{\pi}(s,a)$: $Q^{\pi}(s,a)=\lim_{N\rightarrow\infty}L_{N}(s,a,\pi)$. Although the authors do not consider state-action value functions in their paper, repeating the arguments of asienkiewicz2017note it suffices to consider stationary policies when wanting to solve $\max_{\pi}Q^{\pi}(s,a)$ for all $(s,a)$ and that the solution $Q^{*}$ solves the optimality equation:
$Q^{*}(s,a)=R(s,a)-\frac{\gamma}{\beta}\log\bigg(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime})}]\bigg)\,.$ (8)
@@
Similarly, it is clear that the $Q$-functions satisfy the Bellman recursion relations:
$Q^{\pi}(s,a)=R(s,a)-\frac{\gamma}{\beta}\log\big(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta V^{\pi}(s^{\prime})}]\big)\,.$ (9)
@@ 2.4 Learning Performance
We consider two types of RL algorithms $\mathcal{U}$, namely those that output a $Q$-function $Q_{T}^{\mathcal{U}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ and those that output a policy $\pi_{T}^{\mathcal{U}}:\mathcal{S}\rightarrow\mathcal{A}$ using $T$ transition samples. Note that any algorithm that outputs a $Q$-function also outputs a policy, namely the one obtained by acting greedily with respect to the $Q$-function. There are also ways to obtain a $Q$-function from a policy. There is however no canonical way to do this as the algorithm cannot simply output $Q^{\pi^{\mathcal{U}}}$ since the algorithm does not have access to the true MDP. The way we evaluate the quality of an algorithm that outputs a $Q$-function is by $\|Q^{*}-Q_{T}^{\mathcal{U}}\|$. For an algorithm that instead outputs a policy, we evaluate the policy in terms of $\|V^{*}-V^{\pi_{T}^{\mathcal{U}}}\|$. Often we will suppress $T$ from the notation.
@@
We say that an algorithm $\mathcal{U}$ that outputs a $Q$-function $Q^{\mathcal{U}}$ is $(\varepsilon,\delta)$-correct on a set of MDPs $\mathbb{M}$ if $\mathbb{P}(\|Q^{*}-Q^{\mathcal{U}}\|\leq\varepsilon)\geq 1-\delta$ for all $M\in\mathbb{M}$. Similarly, we say that an algorithm $\mathcal{U}$ that outputs a policy $\pi^{\mathcal{U}}$ is $(\varepsilon,\delta)$-correct on a set of MDPs $\mathbb{M}$ if $\mathbb{P}(\|V^{*}-V^{\mathcal{\pi^{\mathcal{U}}}}\|\leq\varepsilon)\geq 1-\delta$ for all $M\in\mathbb{M}$.
@@ 3 Model-Based Risk-Sensitive Q-Value Iteration
In this section we describe the model-based value iteration algorithm which aims at finding the optimal $Q$-function $Q^{*}$. We then give an upper bound on the total number of calls to the generative model needed in order for this algorithm to be $(\varepsilon,\delta)$-correct. The model based approach is based on working on an MDP, which may disagree with the true MDP because it does not use the true transition probabilities but an estimate of the transition functions obtained from $n$ calls to each of the state-action pairs in $\mathcal{S}\times\mathcal{A}$ as described in the algorithm below.
@@ 12    $\forall s\in\mathcal{S}:\widehat{P}(s,z)=\frac{m(s,z)}{n}$
The model-based approach we describe is general in the sense that any oracle that for any $\varepsilon>0$ can find an $\varepsilon$-optimal policy can be used. We prove the existence of one such oracle in the form of a $Q$-value iteration very like the one from the classical risk-neutral setting. The proof is very similar to Part (a) in Theorem 3.1 in asienkiewicz2017note but is nevertheless provided in Appendix F.
@@
$(\mathcal{T}f)(s,a)$ $=R(s,a)+\frac{-\gamma}{\beta}\log\bigg(\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta\max_{a}f(s^{\prime},a)}\bigg),$
$(\mathcal{T}^{\pi}f)(s,a)$ $=R(s,a)+\frac{-\gamma}{\beta}\log\bigg(\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta f(s^{\prime},\pi(s^{\prime}))}\bigg).$
The operators $\mathcal{T}$ and $\mathcal{T}^{\pi}$ are $\gamma$-contractions with respect to the max-norm, i.e., for value-functions $f_{1}$ and $f_{2}$, it holds that $\|\mathcal{T}f_{1}-\mathcal{T}f_{2}\|\leq\gamma\|f_{1}-f_{2}\|$ and $\|\mathcal{T}^{\pi}f_{1}-\mathcal{T}^{\pi}f_{2}\|\leq\gamma\|f_{1}-f_{2}\|$.
@@ 8        $Q_{j+1}(s,a)=\mathcal{T}Q_{j}(s,a)$
Algorithm 2 RS-QVI(M,k)(M,k)
The next lemma shows that if we choose $k$ large enough in the RS-QVI algorithm, we can obtain $Q_{k}$ and $V^{\pi_{k}}$ that are as close to $Q^{*}$ and $V^{*}$ as we desire. The proof is postponed to Appendix F.
@@
Fix $\varepsilon>0$. Then there exists some $k(\varepsilon)$ such that if the number of iterations in RS-QVI exceeds $k(\varepsilon)$, then the output of Algorithm 2 (RS-QVI) satisfies $\|Q_{k}-Q^{*}\|<\varepsilon$ and $\|V^{\pi_{k}}-V^{*}\|<\varepsilon$.
@@ 3.1 Analysis of MB-RS-QVI
With RS-QVI in place, we need a set of lemmas for the analysis of the sample-complexity of the model based RS-QVI algorithm.
An important result for the analysis is a risk-sensitive version of the simulation lemma kearns2002near ; strehl_analysis_2008 , which describes how different two $Q$-functions for the same policy are in two different MDPs that differ only slightly in their rewards and transition functions. The proof is postponed to Appendix F.
@@
Consider two MDPs $M_{1}=(S,A,P_{1},R,\gamma,\beta)$ and $M_{2}=(S,A,P_{2},R,\gamma,\beta)$ differing only in their transition functions. Fix a stationary policy $\pi$, and let $Q^{\pi}_{1}$ and $Q^{\pi}_{2}$ be respective $Q$-functions of $\pi$ in $M_{1}$ and $M_{2}$. If $\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\|P_{1}(\cdot|s,a)-P_{2}(\cdot|s,a)\|_{1}\leq\tau$. Then, it holds that
$\|Q^{\pi}_{1}-Q^{\pi}_{2}\|\leq\begin{cases}\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-|\beta|(\frac{1}{1-\gamma}-V_{1}(s^{\prime}))}\big|\quad&\beta<0\\ \frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-|\beta|V_{1}(s^{\prime}))}\big|\quad&\beta>0\\ \frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\max_{s,a}\sum_{s^{\prime}\in\mathcal{S}}|P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)|\quad&\beta\neq 0.\end{cases}$
It then follows that if for some $\varepsilon>0$ it holds that any of the $\max_{s,a}$-expressions is smaller than $\varepsilon\frac{1-\gamma}{\gamma}|\beta|e^{-|\beta|\frac{1}{1-\gamma}}$, then $\|Q_{1}^{\pi}-Q_{2}^{\pi}\|\leq\varepsilon$ and ensuring this with high probability is a matter of invoking an appropriate concentration inequality. Using the decompositions from Lemma 4, we get for any $s\in\mathcal{S}$ that
$V^{\pi_{k}}(s)$ $\geq V^{*}(s)-\|V^{\pi_{k}}-\widehat{V}^{\pi_{k}}\|-\|\widehat{V}^{\pi^{*}}-V^{*}\|-\|\widehat{V}^{\pi_{k}}-\widehat{V}^{*}\|,$ (10)
@@
Theorem 1.
There exists a universal constant $c$ such that for any $\varepsilon>0,\delta\in(0,1)$ and any MDP $M$ with $S$ states and $A$ actions, if the learner is allowed to make
$T=c\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|}\bigg)^{2}\log\bigg(\frac{SA}{\delta}\bigg)$ (12)
calls to the generative model, then $\mathbb{P}(\|Q^{*}-Q_{k}\|\leq\varepsilon)\geq 1-\delta$.
@@
For any $\varepsilon>0$, we can get $\|Q_{k}-\widehat{Q}^{*}\|<\varepsilon/2$ using enough iterations of the optimization oracle by Lemma 2. The term $\|\widehat{Q}^{\pi^{*}}-Q^{*}\|$ can also be made smaller than $\varepsilon/2$ by the simulation lemma if either $\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-|\beta|(\frac{1}{1-\gamma}-V_{1}(s^{\prime}))}\big|<\tau$ or $\max_{s,a}\big|\sum_{s^{\prime}\in\mathcal{S}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-|\beta|V_{1}(s^{\prime}))}\big|<\tau$, where $\tau=\frac{\varepsilon}{2}\bigg[\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\bigg]^{-1}$ which can be ensured with probability larger than $1-\delta$ by picking $N=\frac{(1-e^{-\beta\frac{1}{1-\gamma}})^{2}}{2\tau^{2}}\log(2SA/\delta)$. Using that the total calls to the generative model is $T=SAN$ and substituting in the value for $\tau$, we can ensure for all $(s,a)$ that $Q_{k}(s,a)>Q^{*}(s,a)-\varepsilon$ with probability larger than $1-\delta$ by using a total number of samples
$T=3\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|}\bigg)^{2}\log\bigg(\frac{SA}{\delta}\bigg)\,.$ (13)
@@
There exists a universal constant $c$ such that for any $\varepsilon>0,\delta\in(0,1)$, and any MDP $M$ with $S$ states and $A$ actions if the learner is allowed to make
$T=c\frac{SA}{\varepsilon^{2}(1-\gamma)^{4}}\bigg(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|}\bigg)^{2}\log\bigg(\frac{SA}{\delta}\bigg)$ (14)
calls to the generative model then $\mathbb{P}(\|V^{*}-V^{\pi_{k}}\|\leq\varepsilon)\geq 1-\delta$
Proof.
The proof follows immediately from Theorem 1 and Theorem 7 with the choice of $c=6$. ∎
By taking the limit $\beta\rightarrow 0$ leads to PAC-bounds in the risk-neutral case, but the resulting Q-value learning PAC bound is worse by a factor of $\frac{1}{1-\gamma}$ gheshlaghi2013minimax and the policy learning PAC-bound is worse by a factor of $\frac{1}{(1-\gamma)^{3}}$ agarwal_model-based_2020 .
Since the derivations so far lead to a PAC-bound on policy learning that is worse by a factor of $\frac{1}{(1-\gamma)^{2}}$ compared to the PAC bound on Q-value learning, we now provide an alternative derivation that leads to a PAC-bound with the same dependence on horizon. The bound achieved based on the Weissman bound however comes with a worse dependence on $S$, but in cases with long effective horizons and small state-space, this bound might be tighter. Another deficiency of this alternative bound is that it is also less interpretable in the sense that it explodes for $\beta\rightarrow 0$.
Theorem 3.
There exists a universal constant $c$ such that for any $\varepsilon>0,\delta\in(0,1)$, and any MDP $M$ with $S$ states and $A$ actions, if the learner is allowed to make
@@
calls to the generative model, then $\mathbb{P}(\|V^{*}-V^{\pi_{k}}\|\leq\varepsilon)\geq 1-\delta$.
@@
By Lemma 2, we can pick $k$ large enough so that $\|\widehat{V}^{\pi_{k}}-\widehat{V}^{*}\|\leq\varepsilon/3$. By the union bound we have that the two terms $\|V^{\pi_{k}}-\widehat{V}^{\pi_{k}}\|$ and $\|\widehat{V}^{\pi^{*}}-V^{*}\|$ can simultaneously be made smaller than $\varepsilon/3$ by sampling each state-action pair $N=8\frac{S+\log(\frac{2SA}{\delta})}{\tau^{2}}$ times where $\tau=\frac{\varepsilon}{3}\bigg[\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|}\bigg]^{-1}$ such that if we sample
$T=73\bigg[S+\log\bigg(\frac{SA}{\delta}\bigg)\bigg]\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{2|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}$ (16)